OpenStudy (anonymous):

describe how two different ways to use elimination to solve the system {3x-4y=-1 5x+4y=-5

OpenStudy (anonymous):

either u can eliminate x from both equations or y from both equations

OpenStudy (lgbasallote):

i think what she meant is what should you multiply...

OpenStudy (anonymous):

no i think @mayank_mak is right

OpenStudy (anonymous):

@chikitagee so u can solve further.

OpenStudy (anonymous):

@mayank_mak yea but i just dont know how to explain things right like idk how to do this describe how two different ways to use elimination to solve the system {3x-4y=-1 5x+4y=-5

OpenStudy (anonymous):

take the LCM of x coefficient in both equations. In this problem it is 15. so multiply the (3x-4y = -1) with 5 to make 15 and multiply (5x+4y=-5) by 3 to make the coefficient of x 15 then subtract both equations to get value of y

OpenStudy (anonymous):

thanks